# Maud Biquard ISAE-Supaero / CNES

### v1

@@ 5.2 Detailed data processing
Table 2: PCRS (IGN) and Pélican (CNES) databases characteristics. $\downarrow$50cm (resp. 25cm) corresponds to the information of the database downsampled at a resolution of 50cm (resp. 25cm).
@@
In this paper, we use simulated data to train SatDPIR, VBLE-xz, and the deep learning baselines. For the image restoration experiments, both simulated and real satellite data are used. The simulated images are produced using two databases: PCRS, provided by IGN [1], and Pélican, provided by CNES. Both datasets are composed of airplane images acquired at a very high resolution, that is 5cm for PCRS and 10cm for Pélican. Target and degraded satellite images are then obtained by downsampling them. The characteristics of each database are provided in Tab. 2.
5.2.1 Simulated satellite images
@@
To train DR methods SatDPIR and VBLE-xz, as well as direct inversion baselines, target satellite images are required. Simulated degraded images are also needed to evaluate the performance of the different methods, as well as to train direct inversion baselines. To create target images, very high resolution PCRS and Pélican images are downsampled at the target resolution, that is 50cm for IR and 25cm for IR+SISR, after applying an anti-aliasing filter. By downsampling these images by a sufficient factor, we consider the initial airplane blur to be negligible, thus creating ”ideal” noise and blur free satellite images. To create the degraded images, the blur kernel, modelling the satellite point spread function (PSF), is applied to Pélican and PCRS aerial images, before downsampling them at the satellite resolution, that is 50cm in all simulated experiments, adding the instrument noise and quantizing them on 12 bits. Note that in this simulation pipeline, degraded images do not exactly follow our forward model of eq. 18, because of PSF subsampling and image quantization, which is coherent with a realistic context where the forward model is always imperfect.
5.2.2 Real satellite images
@@
Experiments on real data are performed using Pléiades 12 bits images. These images are acquired at a resolution of 70cm (but commercialized at 50cm after an upsampling that enhances the image robustness to post-processing). Note that, unlike the simulated images, the real images have been compressed at 2.86bits/pixel [40] and decompressed. For these images, we choose to neglect the compression when modelling the data log-likelihood, which is required for SatDPIR and VBLE computations. This is relevant as the compression rates for satellite images are very high. Therefore, we consider the impact of compression-decompression as a source of error, among others, in the real image model.
5.2.3 Train and test datasets
For the simulated test set, a subset of 30 images from the Pélican dataset is chosen, representing various landscapes such as desert, industrial, and urban areas. Degraded images are simulated at a resolution of 50cm, and are sized $820\times 820$. We denote this test set as TEST30_SIMU. All PCRS images and the other Pélican images are used in the training and validation set. For some experiments, only a subset of TEST30_SIMU is used, denoted TEST14_SIMU. It contains 14 images, center cropped to sizes $512\times 512$ for IR and $256\times 256$ for IR+SISR experiments. To test the algorithms on real Pléiades images, three landscapes are chosen: ATACAMA, DENVER, and SHANGHAI to form TEST30_REAL dataset. It contains 10 images of each landscape, each image being $820\times 820$ at a resolution of 70cm.
@@ 5.3 Experiment setup
5.3.1 General training settings
The neural networks used in VBLE-xz and SatDPIR are trained at each used target resolution, that is 50cm and 25cm for the IR and IR+SISR experiments on simulated data, 70cm and 35cm for the IR and IR+SISR experiments on real data.
5.3.2 Compressive autoencoder details
For VBLE-xz, we choose CAE architecture from [52]. It is a CAE with hyperprior, with a simple convolutional encoder and decoder. We employ models from compressAI library [8] that have been pretrained at different bitrates and we finetune them on satellite images at a given target resolution for 150k iterations. Afterward, we train the variance decoder as described in Sec. 4.5 using satellite images during 100k iterations.
5.3.3 Denoiser details
DRUNet denoiser [76] is used for SatDPIR. DRUNet is a state-of-the-art deep convolutional denoiser specifically tailored for image restoration. We employ a pretrained model from [76] that we finetune at a given target resolution on satellite images, during 500k iterations.
SRResNet RDN DRUNet CAE
#params 1.3M 22M 32M 25M
Table 3: Number of parameters of the different networks used in the experiments
@@ 5.3.4 Baselines
SatDPIR and VBLE-xz are compared to several baselines to assess their performance. First, for uncertainty quantification evaluation, we compare VBLE-xz to two posterior sampling methods: PnP-ULA [41], a state-of-the-art MCMC posterior sampling method that employs a denoiser as prior; and VBLE [5], which is a natural baseline for VBLE-xz. For PnP-ULA, we choose the DRUNet denoiser, that is the denoiser used for SatDPIR, and for VBLE we choose the same CAE as for VBLE-xz. Then, to assess the performance of VBLE-xz and SatDPIR without uncertainty quantification, we consider the following baselines. Bay+IF [40] is a classical approach that is currently being used in the Pléiades ground segment. It is composed of two steps: denoising is performed using NL-Bayes [42], then inverse filtering is used for deblurring. For super-resolution experiments, Bay+IF images are upsampled using bicubic upsampling. Then, we consider two direct inversion baselines, SRResNet and RDN. SRResNet [43] is a light convolutional neural network composed of a succession of residual blocks. This network has already been adapted for satellite image restoration and SISR [81, 12]. RDN [79] is a deeper and wider convolutional network, which uses residual blocks with dense connections. For both networks, we adopt the loss proposed in [12]: $\mathcal{L}(x,\hat{x})=\alpha\text{SmoothL1}(x,\hat{x})+\beta\operatorname{DISTS}(x,\hat{x})$ where SmoothL1 is a smooth version of the L1 distance, and DISTS [16] is a perceptual loss. SRResNet and RDN are trained from scratch during 200k iterations, corresponding to convergence. In Tab. 3, the number of parameters of every neural network used in the experiments is provided.
@@ 5.3.5 Restoration hyperparameters
For several methods, restoration hyperparameters are tuned by grid search on a validation set of 3 Pélican images for each inverse problem. For SatDPIR, the regularization parameter $\lambda$ is tuned accordingly. For VBLE-xz (and VBLE), we also introduce a regularization parameter $\lambda$ in eq. 15 as follows:
$\arg\max_{\bar{z},a,b}\mathbb{E}_{q_{b}(x|z)q_{\bar{z},a}(z)}\big{[}\log p_{Y|X}(y|x)+\lambda\big{(}\log p_{\theta}(x|z)$
$+\log p_{\theta}(z)-\log q_{b}(x|z)-\log q_{\bar{z},a}(z)\big{)}\big{]},$ (19)
$\lambda=1$ corresponding to the Bayesian framework. We choose $\lambda=0.6$ in all VBLE and VBLE-xz experiments because it yields the best trade-off between performance and uncertainties quality, as will be shown in Fig. 12 of the ablation study. For PnP-ULA, the regularization parameter, the denoiser noise level, and the chain step size are tuned accordingly, while $10^{5}$ iterations are performed for each restoration.
@@ 5.3.6 Evaluation metrics
To evaluate the image restoration performance when ground truth is available, three metrics have been used: the Peak Signal-to-Noise Ratio (PSNR), the Structural SIMilarity (SSIM) [67], and the Learned Perceptual Image Patch Similarity (LPIPS) [78]. The PSNR represents the accuracy between the restored and target images, while the SSIM and LPIPS are respectively classical and deep learning perceptual metrics. For no reference image quality assessment, we employ the Blind Referenceless Image Spatial Quality Evaluator (BRISQUE) [53]. BRISQUE provides a score between 0 and 100, based on the deviation of the statistic of some local features with respect to natural images, the lower being the better. Finally, to evaluate quantitatively the predicted uncertainty quantification of posterior sampling methods, we compute Interval Coverage Probability (ICP) of level $\alpha$. It represents the proportion of the perpixel true error that is contained in the perpixel predicted error quantiles of level $\alpha$. Thus, the closer to $\alpha$, the better.
@@ 6 Experimental results
6.1 Uncertainty quantification results on simulated data
IR PnP-ULA VBLE
VBLE-xz
not calibrated
VBLE-xz
PSNR $\uparrow$ 48.63 48.40 48.35 48.35
ICP $90\%$ 0.92 0.79 0.83 0.91
Time 2h10m 17.7sec 27.3sec 27.3sec
IR+SISR PnP-ULA VBLE
VBLE-xz
not calibrated
VBLE-xz
PSNR $\uparrow$ 35.86 37.45 37.38 37.38
ICP $90\%$ 0.94 0.67 0.73 0.91
Time 2h21m 19.5sec 29.4sec 29.4sec
@@ 6.1 Uncertainty quantification results on simulated data
Table 4: Comparison between posterior sampling methods on TEST14_SIMU for IR and IR+SISR. Time denotes the time required to restore and then draw 100 posterior samples for each method given a $512^{2}$ image for IR and $256^{2}$ for IR+SISR. Times are obtained with a Nvidia Quadro RTX8000 GPU.
@@
Figure 3: Interval coverage probabilities for the IR problem (left) and IR+SISR problem (right). Each point of a curve of abscissa $\alpha$ provides the ICP of level $\alpha$ averaged on the test dataset. Colored areas correspond to deviations. The closer to the identity, the better.
Figure 4: Difference between VBLE and VBLE-xz posterior sampling ability. The true error is $e=|GT-MMSE|$ while the error quantile map represents the perpixel 90% quantile of the predicted error. The calibrated version refers to the improvement proposed in Sec. 4.6. ©CNES 2024
@@
Figure 5: Posterior sampling ability of various methods on TEST14_SIMU. Top: IR problem, Bottom: IR+SISR problem. MMSE estimates are averaged on 100 posterior samples. The true error is $e=|GT-MMSE|$ while the error quantile map represents the perpixel 90% quantile of the predicted error. VBLE-xz quantiles are calibrated, as proposed in Sec. 4.6. ©CNES 2024
@@
In this section, we assess the performance and interest of the posterior samples and uncertainty quantification provided by VBLE-xz compared to VBLE [5] and PnP-ULA [41]. We perform these experiments on TEST14_SIMU. In Tab. 4 we report the PSNR, the 90% ICP (defined in Sec. 5.3.6) as well as the computation time required to restore an image and draw 100 posterior samples, with best values in bold font. First, a key aspect of VBLE-xz is its computation time to draw 100 posterior samples. Indeed, it is by several orders of magnitude faster than PnP-ULA, making this method scalable for satellite image restoration, unlike PnP-ULA. In terms of PSNR, PnP-ULA exhibits consistent results for image restoration, but it is outperformed by VBLE and VBLE-xz in super-resolution. In Fig. 3, coverage probabilities, that is ICP of level $\alpha$ for $\alpha$ varying from 0 to 1, are provided for each method on IR and IR+SISR. The 90% ICP of Tab. 4 corresponds to the point with abscissa 0.9. First, we observe that VBLE-xz without calibration obtains better 90% ICPs than VBLE, showing the interest of modelling the reconstruction error. While empirical coverage probabilities show that VBLE-xz without calibration still underestimates the uncertainty, the calibration procedure solves the problem, enabling our method to be competitive with PnP-ULA ICPs for IR, and clearly outperforming it for IR+SISR.
@@
Additionally, the interest of VBLE-xz over VBLE is visually assessed in Fig. 4. This figure shows the true perpixel error, as well as the 90% perpixel predicted error quantiles for VBLE, VBLE-xz without calibration, and VBLE-xz with calibration. VBLE-xz, even without calibration, possesses more realistic quantiles, as the quantile map has higher values and exhibits higher uncertainties on some uniform zones as roofs, while VBLE seems to concentrate the uncertainty on the edges. With calibration, VBLE-xz is able to provide very accurate uncertainty quantification, as the quantile map resembles the true perpixel error. Lastly, a thorough visual study of each method’s posterior sampling ability is provided in Fig. 5. This figure shows the MMSE estimate of each method obtained by averaging 100 posterior samples, as well as two posterior samples, the true error, and the 90% error quantile map. PnP-ULA posterior samples are very noisy for IR+SISR, making them hardly usable in practice. VBLE samples are visually better than those of PnP-ULA, but some details are unrealistic. In particular, in the shadow, the cars are sometimes distorted. VBLE-xz samples exhibit a little bit of noise as some uncorrelated noise is added in the image space, but this does not affect the image interpretation and VBLE-xz gets more realistic samples than VBLE. At last, the 90% error quantiles of VBLE-xz are consistent with the true error.
@@
Hence, VBLE-xz provides better posterior samples and uncertainties than VBLE, validating the proposed joint latent and image posterior modelling. Furthermore, VBLE-xz outperforms PnP-ULA in terms of sampling ability, and computation speed. In particular, its low computation time for a posterior sampling method makes it scalable to restore real satellite images.
@@ 6.2 Point estimation results on simulated data
Bay+IF RDN SRResNet SatDPIR VBLE-xz
@@
RDN SRResNet SatDPIR VBLE-xz
DENSE URBAN PSNR $\uparrow$ 30.74 38.12 47.00 47.86 47.89 47.42 31.14 35.27 34.89 35.61 35.25
SSIM $\uparrow$ 0.9083 0.9841 0.9964 0.9964 0.9964 0.9958 0.8929 0.9453 0.9415 0.9491 0.9447
LPIPS $\downarrow$ 0.1477 0.0233 0.0065 0.0067 0.0063 0.0104 0.2366 0.1028 0.1219 0.1450 0.1618
LIGHT URBAN PSNR $\uparrow$ 31.16 38.59 47.50 47.98 48.09 47.61 31.49 36.00 35.52 36.42 35.83
SSIM $\uparrow$ 0.9155 0.9849 0.9960 0.9958 0.9960 0.9954 0.9037 0.9512 0.9474 0.9554 0.9506
LPIPS $\downarrow$ 0.1436 0.0280 0.0089 0.0095 0.0085 0.0153 0.2416 0.1067 0.1280 0.1478 0.1669
RURAL PSNR $\uparrow$ 36.30 42.60 48.01 48.04 48.20 47.70 35.27 37.45 37.28 38.05 37.47
SSIM $\uparrow$ 0.9346 0.9830 0.9918 0.9917 0.9919 0.9907 0.8972 0.9232 0.9229 0.9319 0.9211
LPIPS $\downarrow$ 0.1722 0.1045 0.0263 0.0273 0.0255 0.0484 0.3060 0.1729 0.1942 0.2338 0.2831
DESERT PSNR $\uparrow$ 44.06 47.04 48.49 48.57 48.74 47.89 40.01 39.65 39.87 40.60 40.05
SSIM $\uparrow$ 0.9703 0.9830 0.9876 0.9875 0.9878 0.9850 0.9154 0.9150 0.9126 0.9256 0.9128
LPIPS $\downarrow$ 0.2058 0.1045 0.0511 0.0513 0.0514 0.1181 0.4052 0.2263 0.2632 0.3212 0.3764
WATER PSNR $\uparrow$ 36.54 44.01 52.84 51.86 53.07 52.82 36.09 40.01 39.38 40.44 40.12
SSIM $\uparrow$ 0.9705 0.9945 0.9984 0.9977 0.9983 0.9982 0.9650 0.9810 0.9785 0.9829 0.9810
LPIPS $\downarrow$ 0.0877 0.0302 0.0101 0.0178 0.0091 0.0193 0.1075 0.0658 0.1133 0.0758 0.0801
PORT PSNR $\uparrow$ 30.28 37.91 46.96 48.03 48.20 47.86 29.38 32.73 32.31 32.97 32.47
SSIM $\uparrow$ 0.9401 0.9897 0.9974 0.9972 0.9975 0.9971 0.9152 0.9517 0.9474 0.9549 0.9496
LPIPS $\downarrow$ 0.1065 0.0313 0.0108 0.0129 0.0089 0.0225 0.1785 0.0945 0.1263 0.1175 0.1345
FOREST PSNR $\uparrow$ 31.53 38.87 47.28 47.30 47.39 47.12 32.00 34.72 34.51 35.10 34.93
SSIM $\uparrow$ 0.9184 0.9862 0.9965 0.9964 0.9965 0.9961 0.9048 0.9429 0.9410 0.9485 0.9450
LPIPS $\downarrow$ 0.1689 0.0334 0.0115 0.0124 0.0109 0.0185 0.2567 0.1240 0.1472 0.1938 0.2248
INDUSTRIAL PSNR $\uparrow$ 35.52 42.45 48.59 48.99 49.29 48.73 34.89 38.17 37.69 38.65 38.33
SSIM $\uparrow$ 0.9358 0.9863 0.9943 0.9940 0.9944 0.9935 0.9143 0.9481 0.9447 0.9543 0.9469
LPIPS $\downarrow$ 0.1352 0.0398 0.0169 0.0179 0.0160 0.0333 0.2573 0.1339 0.1542 0.1693 0.1894
TOTAL PSNR $\uparrow$ 33.55 40.56 48.08 48.44 48.66 48.19 33.22 36.66 36.28 37.18 36.63
SSIM $\uparrow$ 0.9289 0.9859 0.9951 0.9949 0.9952 0.9944 0.9088 0.9460 0.9430 0.9513 0.9452
LPIPS $\downarrow$ 0.1437 0.0369 0.0145 0.0157 0.0138 0.0275 0.2463 0.1228 0.1466 0.1658 0.1897
Computation time x x 4.2s 0.4s 3.7s 62.3s x 5.0s 0.9s 1min08s 3min20s
@@
Figure 6: Qualitative results of SatDPIR, VBLE-xz and the baselines on TEST30_SIMU. The first two lines correspond to the IR problem, the others correspond to IR+SISR. ©CNES 2024
In this section, we present the results of the different methods on the TEST30_SIMU dataset. Quantitative metrics are presented in Tab. 5 while Fig. 6 contains visual results. Table 5 contains results for the image restoration problem, with (IR+SISR) and without super-resolution (IR). The results have been separated for each type of landscape, enabling a thorough comparison between the methods. First, the Bay+IF method is by far outperformed by all deep learning methods. Then, SatDPIR outperforms direct inversion methods RDN and SRResNet in terms of PSNR and SSIM, while being slightly below them in terms of LPIPS. This is expected because inversion methods are partly trained with a perceptual loss, DISTS, which is similar to LPIPS, whereas DR methods only consider the pixelwise likelihood. This can be visually illustrated in Fig. 6. Indeed, direct inversion methods exhibit sharper results with more high frequency details, which is favorable to a low LPIPS. However, the added high frequencies are not necessarily correct. Moreover, they show non desirable oscillating deconvolution artifacts, for example in the 5th image of Fig. 6. Compared to the other methods, VBLE-xz exhibits slightly poorer metrics than SatDPIR, but remains competitive with the baselines in terms of PSNR and LPIPS. Visually, VBLE-xz is particularly interesting as its MMSE result is a bit smoother than the other method, but exhibits very few artefacts, compared to direct inversion methods, and also to SatDPIR, which invents some lines in the 3rd image of Fig. 6.
@@
The computation time needed to restore one image is also provided in Tab. 5, except for Bay+IF, which runs on CPU. First, the computation time is much higher when super-resolution is performed, as it is necessary to divide images into patches for all methods except SRResNet. Then, the computation time of SatDPIR without super-resolution is very fast, and even faster than RDN. When super-resolution is performed, direct inversion methods are faster, even if DR methods’ computation time remains scalable. It is worth noting that SatDPIR computation time has been greatly improved compared to DPIR by our proposed adaptation in Sec. 3, as will be detailed in Tab. 8 of the ablation study. Concerning VBLE-xz, its computation time is slightly above the others. This is expected as VBLE-xz enables posterior sampling. Yet, it remarkably stays computationally scalable. Besides, recall that this is rarely the case with other posterior sampling methods as VBLE-xz is much faster than PnP-ULA as depicted in Tab. 4 of the previous section.
@@
Therefore, SatDPIR provides excellent results as it reaches direct inversion methods performances, and even outperforms them in terms of PSNR, even though SatDPIR is a plug-and-play method that has not been trained to solve a specific inverse problem. It is also very fast for IR, a bit slower than RDN and SRResNet for IR+SISR but stays computationally efficient. Furthermore, VBLE-xz, whose posterior sampling ability has been demonstrated in the previous section, exhibits very consistent results compared to the baselines, which only compute a point estimate. Its computation time is slightly bigger, however stays remarkably low for a posterior sampling method.

### v2

@@ 4.2 Detailed data description
DATASET #imgs $y$ size resolution
TEST30 30 $820\times 820$ 50cm
TEST14_PS512 14 $512\times 512$ 50cm
TEST14_PS256 14 $256\times 256$ 50cm
TEST30_REAL 30 $820\times 820$ 70cm
@@
Table 1: Top: PCRS (IGN) and Pélican (CNES) databases characteristics. $\downarrow$50cm (resp. 25cm) corresponds to the information of the database downsampled at a resolution of 50cm (resp. 25cm). Bottom: Characteristics of the test databases used in the experiments. TEST30_REAL contains real satellite images, the other images are simulated.
4.2.1 Test and train datasets
@@
In this paper, we use simulated data to train VBLE-xz and the baselines. For the image restoration experiments, both simulated and real satellite data are used. The simulated images are produced using two databases: PCRS, provided by Institut national de l’information géographique et forestière (IGN) [IGNPCRS], and Pélican, provided by CNES. Both datasets are composed of airplane images acquired at a very high resolution, that is 5cm for PCRS and 10cm for Pélican. Target and degraded satellite images are then obtained by downsampling them. The characteristics of each database are provided in the first part of Tab. 1. For the simulated test set, a subset of 30 images from the Pélican dataset is chosen, representing various landscapes such as desert, industrial, and urban areas. All PCRS images and the other Pélican images are used in the training and validation set. Some tests are performed on subsets of this test set because of the high computational cost of some baselines. The various test subsets are detailed in the second part of Tab. 1. The tests on real Pléiades images are performed on 30 of them. Their characteristics are also given in Tab. 1.
4.2.2 Image simulation process
@@
Target images are obtained by downsampling very high-resolution PCRS and Pélican images at the target resolution, that is 50cm for IR and for inpainting, and 25cm for IR+SISR, after anti-aliasing filtering. A sufficiently high downsampling factor makes the initial airplane degradation negligible, thus allowing us to consider them as ideal, that is free from noise and blur. Degraded images for IR and IR+SISR are obtained by first applying the blur kernel, which models the satellite point spread function (PSF), to Pélican and PCRS aerial images. Second, they are downsampled at the satellite resolution, that is 50cm in all simulated experiments, subjected to the simulated instrument noise, and finally quantized on 12 bits. Note that degraded images obtained with this simulation pipeline do not strictly follow the forward model eq. 15 that does not account for PSF subsampling and image quantization. This places us in a realistic context where the forward model is imperfectly known.
4.2.3 Characteristics of the real images
@@
Experiments on real data are performed using Pléiades 12 bits images. These images are acquired at a resolution of 70cm (but commercialized at 50cm after an upsampling that enhances the image robustness to post-processing). Note that, unlike the simulated images, the real ones have been compressed at 2.86bits/pixel [Latry2012] and decompressed. At such a high rate, the impact of compression-decompression can also be considered as an admissible practical deviation from the theoretical model.
@@ 4.3 Experiment setup
4.3.1 Training settings
The neural networks used in VBLE-xz are trained at each considered target resolution, that is 50cm and 25cm for the IR and IR+SISR experiments on simulated data, 70cm and 35cm for the IR and IR+SISR experiments on real data. We choose CAE architecture from [Minnen2018]. It is a CAE with a hyperprior, with a simple convolutional encoder and decoder. We employ a pretrained model from compressAI library [Begaint2020] that has been pretrained at a high bitrate and we finetune it on satellite images at a given target resolution for 150k iterations. Afterward, we train the variance decoder as described in Sec. 3.4.3 using satellite images during 100k iterations.
SRResNet RDN DRUNet DiffUNet CAE
#params 1.3M 22M 32M 552M 25M
Table 2: Number of parameters of the different networks used in the experiments. DiffUnet corresponds to the diffusion model used in DiffPIR.
@@ 4.3.2 Baselines
[…]
@@ 4.3.3 Restoration hyperparameters
For several methods, restoration hyperparameters are tuned by grid search on a validation set of 3 Pélican images for each inverse problem. This procedure was applied to: the regularization parameter $\lambda$ and the scaling $s$ for VBLE-xz, the regularization parameter for PG-DPIR, the regularization parameter, the denoiser noise level and the chain step size for PnP-ULA and the regularization parameter for DiffPIR. For PnP-ULA, $10^{5}$ iterations are performed for each restoration. For DiffPIR, 100 diffusion steps are performed and $\zeta$ is fixed to 0.9.
@@ 4.3.4 Evaluation metrics
To evaluate the image restoration performance when ground truth is available, three metrics have been used: the Peak Signal-to-Noise Ratio (PSNR), the Structural SIMilarity (SSIM) [Wang2004], and the Learned Perceptual Image Patch Similarity (LPIPS) [Zhang2018]. The PSNR represents the accuracy between the restored and target images, while the SSIM and LPIPS are respectively classical and deep learning perceptual metrics. For no reference image quality assessment, we employ the Blind Referenceless Image Spatial Quality Evaluator (BRISQUE) [Mittal2012]. BRISQUE provides a score between 0 and 100, based on the deviation of the statistic of some local features with respect to natural images, the lower being the better. Finally, to evaluate quantitatively the predicted uncertainty quantification of posterior sampling methods, we compute Interval Coverage Probability (ICP) of level $\alpha$. It represents the proportion of the perpixel true error that is contained in the perpixel predicted error quantiles of level $\alpha$. Thus, the closer to $\alpha$, the better.
@@ 5 Experimental results
5.1 Comparison between posterior sampling methods on simulated data
PnP-ULA VBLE DiffPIR DiffPIR-100 VBLE-xz
[…]
Time 2h10m 23.5s x 1h55min 29.5s
[…]
LPIPS $\downarrow$ 0.2272 0.1773 0.1746 0.1668 0.1691
[…]
ICP $90\%$ 0.47 0.67 x 0.79 0.90
Time 2h10min 33.3s x 1h57min 43.9s
@@ 5.1 Comparison between posterior sampling methods on simulated data
Table 3: Comparison between posterior sampling methods on TEST14_SIMU for IR, IR+SISR and inpainting. Time denotes the time required to restore and draw 100 posterior samples for each method given a $512^{2}$ image for IR and $256^{2}$ for IR+SISR. For VBLE-xz, it includes the restoration time of algorithm 1 and the posterior sampling time. Times are obtained with a Nvidia Quadro RTX8000 GPU.
@@
Figure 4: Interval coverage probabilities for the IR problem (left), IR+SISR problem (middle) and inpainting (right). Each point of a curve of abscissa $\alpha$ provides the ICP of level $\alpha$ averaged on the test dataset. Colored areas correspond to deviations. The closer to the identity, the better.
@@
Figure 5: Posterior sampling ability of various methods on TEST14_SIMU. Top: Structured inpainting problem. Bottom: IR+SISR problem. MMSE estimates are averaged on 100 posterior samples. The true error is e=|GTMMSE|e=|GT-MMSE| while the error quantile map represents the perpixel 90% quantile of the predicted error. ©CNES 2025
@@
In this section, we assess the performance in restoration and for uncertainty quantification, for the posterior sampling methods. VBLE-xz is thus compared to DiffPIR [Zhu2023]VBLE [Biquard2023] and PnP-ULA [Laumont2022a]. We perform these experiments on TEST14 sets, for IR, IR+SISR and inpainting. We provide quantitative results in Tab. 3, including the 90% ICP (defined in Sec. 4.3.4) as well as the computation time required to restore an image and draw 100 posterior samples. For DiffPIR, the results are given using one posterior sample, as classically done in the literature, and with the posterior mean averaged on 100 samples, denoted as DiffPIR100. First, a key advantage of VBLE-xz is the little time it takes to draw 100 posterior samples. Indeed, it is by several orders of magnitude faster than PnP-ULA and DiffPIR100, making this method scalable for satellite image restoration. In terms of PSNR and LPIPS, VBLE and VBLE-xz have very close results, with VBLE-xz being slightly better, especially in inpainting. This gain is not significant but is a positive outcome as VBLE-xz was primarily designed to enhance uncertainty quantification in VBLE. PnP-ULA exhibits consistent results but it is outperformed by VBLE-xz in super-resolution. DiffPIR is outperformed by VBLE-xz in all cases, while DiffPIR100 performs very well, however with a significant computation time, making it unrealistic for real satellite processing chains. In Fig. 4, coverage probabilities, that is ICP of level $\alpha$ for $\alpha$ varying from 0 to 1, are provided for each method on the three problems. The 90% ICP of Tab. 3 corresponds to the point with abscissa 0.9. First, VBLE-xz yields significantly better uncertainties than VBLE, validating the proposed joint latent and image posterior modeling. Besides, on average, VBLE-xz outperforms PnP-ULA and is close to DiffPIR100 in terms of ICP.
@@
Lastly, a thorough visual study of each method’s posterior sampling ability is provided in Fig. 5. This figure shows the MMSE estimate of each method obtained by averaging 100 posterior samples, as well as two posterior samples, the true error, and the 90% error quantile map. The interest of VBLE-xz over VBLE is visually assessed in the inpainting problem, as the inpainting mask is not retrieved in VBLE predicted error. Indeed, this kind of uncertainty, which presents a very sharp structure, is hardly modeled in the latent space only. We note that VBLE-xz posterior samples exhibit a little bit of noise as some uncorrelated noise is added in the image space, but this does not affect the image interpretation. In contrast, in the particular case of IR+SISR, PnP-ULA posterior samples are very noisy making them hardly usable in practice, while DiffPIR100 and VBLE-xz yield realistic posterior samples and predicted errors.
@@
Hence, VBLE-xz provides better posterior samples and uncertainties than VBLE, validating the proposed joint latent and image posterior modeling. Furthermore, VBLE-xz is faster than DiffPIR100 and PnP-ULA by several orders of magnitude for posterior sampling. This makes it scalable to restore real satellite images, unlike those methods.
@@ 5.2 Point estimation results on simulated data
Bay+IF RDN SRResNet DiffPIR VBLE PG-DPIR VBLE-xz
@@
RDN SRResNet DiffPIR VBLE PG-DPIR VBLE-xz
[…]
Time x x 4.2s 0.4s 3min34s 49.8s 3.7s 1min01s x 5.0s 0.9s 18min42s 4min05s 1min08s 5min10s
@@
Figure 6: Qualitative results of VBLE-xz and the baselines on TEST30_SIMU. The first two images correspond to the IR problem, the others correspond to IR+SISR. The error map is the absolute error between the image and GT (Ground Truth).©CNES 2025
In this section, we present the results of different methods on the TEST30 dataset. In particular, we aim to compare VBLE-xz and PG-DPIR, which are deep regularization (DR) methods, to direct inversion methods RDN and SRResNet, which are mainly chosen in satellite image processing chains. Quantitative metrics are presented in Tab. 4 while Fig. 6 contains visual results. Table 4 contains results for the image restoration problem, with (IR+SISR) and without super-resolution (IR). First, the Bay+IF method is by far outperformed by all deep learning methods. VBLE-xz and PG-DPIR outperform DiffPIR and are competitive, or above direct inversion in terms of PSNR and SSIM, while being slightly below them in terms of LPIPS. This is expected because inversion methods are partly trained with a perceptual loss, DISTS, which is similar to LPIPS, whereas DR methods only consider pixelwise likelihood. Interestingly, in the IR task, SRResNet achieves a higher PSNR than RDN, despite RDN being deeper and theoretically expected to perform better. As the test set differs from the trained images, it is likely that SRResNet generalizes better than RDN, which is plausible given that SRResNet is much shallower than RDN.
Visual results are available in Fig. 6. Direct inversion methods exhibit sharper results with more high-frequency details, which is favorable to a low LPIPS. However, the added high frequencies are not necessarily correct. Moreover, they show non-desirable oscillating deconvolution artifacts, for example in the 4th image of Fig. 6. VBLE-xz is particularly interesting because its MMSE result is slightly smoother than the estimates of the other methods, but exhibits very few artifacts compared to direct inversion methods, as well as to PG-DPIR, which, for instance, invents some lines in the third image of Fig. 6.
@@
The computation time needed to restore one image is also provided in Tab. 4, except for Bay+IF, which runs on CPU. First, the computation time is much higher when super-resolution is performed, as it is necessary to divide images into patches for all methods except SRResNet. Then, direct inversion are faster than DR methods VBLE-xz and PG-DPIR even if DR method’s computation time remains scalable. Concerning VBLE-xz, its computation time is better than DiffPIR one, yet remains above direct inversion computation times. This is expected as VBLE-xz enables posterior sampling. Yet, it remarkably stays computationally scalable. Besides, recall that this is rarely the case with other posterior sampling methods as depicted in Tab. 3 of the previous section.
@@
Therefore, PG-DPIR and VBLE-xz provide excellent results, similar to direct inversion methods in terms of performance. This is remarkable as these methods have not been trained to solve a specific inverse problem. Besides, PG-DPIR is very fast for IR, a bit slower than RDN and SRResNet for IR+SISR but remaining computationally efficient. VBLE-xz is slightly slower, however, it stays remarkably fast for a posterior sampling method.